1. You have to make a 90,000,000 payment in Japanese Yen on close of business day, Friday, January 18th. You decide to hedge your risk with the futures contracts. Assume you that you enter into the futures position at a close of day on Tuesday, January 15th. Futures and spot data are provided in the file BELOW. Contract size is 12,500,000 yen.
    1. Describe the position you decide to enter (long or short).
    1. Describe the contract (what month, and what quantity).
    1. Document the gain or loss due to marking to market every day that your position is open.
    1. What is the total cost in US$ after you have closed out your futures positions, and made your payment?
    1. What would have been the total cost in US$, if you had not hedged? Did you benefit from hedging?
    1. Repeat steps a) – e) assuming that you expect to receive a payment of 900,000 British pounds on close of business day, Friday, January 18th.  Assume you that you enter into the futures position at a close of day on Tuesday, January 15th. Futures and spot data are provided in the file BELOW. Contract size is 62,500 British pounds.

DATA

Japanese Yen Data Daily Settlements for Japanese Yen Future (FINAL Trade Date: Wednesday, 01/16/2019 Estimated Prior Day Month Open HighLowLast Change |Settle FEB 19 92455 92500 91850 91850 -290 92060 525 MAR 19 92480 92730 92025 92115B-29092265 | 103,346 APR 19 92760 92860B 92340A 92340A -290 92500 1,745 MAY 19 JUN 19 93300 193345B 92750 92750295 192970 SEP 19 Volume lOpen Interest 787 225,891 1,451 28 901 110 93080B 92560A 92560A -290 92710 0 36 300 93690 Daily Settlements for Japanese Yen Futures (FINAL Trade Date: Fridav, 01/18/2019 EstimatedPrior Day Month [Open HighLowLast Change |Settle FEB 19 91730 91820B 91240 91320B-44591300 MAR 19 91980 92115 9140591500-44591500 111,197 APR 19 |91930 92240B 91720A 91740B-445 91730 MAY 19 JUN 19 192595 92700B 92190A 92225-45092200 SEP 19 Volume Open Interest 403 675 217,141 563 148 1,012 110 358 92460B 91940A 91960B450 [91940 169 -440 92920 Spot data: YYYY/MM/DD|wdyl USD/JPY 2458499 2019/01/15 Tue o.0092115 2458500 2019/01/16 Wedo.0091863 2458501 2019/01/17Thu 0.0091771 2458502 2019/01/18 Frio.0091176
Answer #1

Ans.

a) Since payment is needed to be made in Japanese Yen, hence to avoid exchange risk one is needed to buy Japanese Yen Futures contract.

b) As the payment is required to deliver in January itself and most Liquid Contract available is MARCH 19 (liquid contract has lesser Impact cost) so we should choose MARCH19 contract to hedge the position.

Position Size: Amount to be hedged/ Contract Size

= 90,000,000/12,500,000

= 7.2 contract needed to buy

But since contract size is standardized hence fractions are not allowed so we should buy 7 contract to hedge our positions.

c) Mark to Market (M2M Position) :

For one point movement in JPY value of contract move by $1250

Contract is entered at price of 92.265+0.29 =92.555

on 16th Jan JPY fallen by 0.29 points settled at 92.265 so M2M will be 7*1250*0.29=$2537.5 minus (loss on positions)

on 17th Jan JPY fallen by 0.32 points settled at 91.945 so M2m will be 7*1250*.32= $2800 minus (loss positions)

and for 18th Jan fallen by 0.445 points so M2M will be 7*1250*0.445=$3893.75 minus (loss on positions)

d) Hence the total cost incurred will be the loss/profit on the positions

as mentioned above the M2M minus = $(2537.5+2800+3893.75)=$9231.3 (loss)

e) If the position would not have been hedged then one would have been stand to gain 0.939 on per JPY and by hedging it actually incurred loss in comparison to without hedged position and situation may have just oppose in case if JPY moved up during the same time.